Hi Can I have some help with this equation?
Find an equation of the line through (-5, 6) with slope -10.

Answer by checkley71(8403)

Y=mX+b
6=-10*-5+b
6=50+b
b=-50+6
b=-44
THUS THE EQUATION IS:
Y=-10X-44